OpenStudy (anonymous):

sole by elimination 9x+6y=-30 4x-4y=20

OpenStudy (anonymous):

representing 9x+6y=-30 as equation one and 4x-4y=20 by equation two. multiply equation one by 2 and equation two by 3. you will get 18x+12y=-60 and 12x-12y=-60 adding the two new equations, 6x=0 therefore x=0. subtituting x=0 into equation one or two and solving it gives y=-5

OpenStudy (anonymous):

oh ok thx